Question 1

Solve the following equation

a) 2e

2x−1 + 5e

x

2

= 0

b) 2

4x + 22x−1 > 2

[8,6]


Question 2

Let f : (−∞, 2] → R be given by

f(x) = √

2 − x

a) Show that f is injective.

b) Determine im(f).

c) Find a left inverse g : R → (−∞, 2] of f.

[7,6,5]


Question 3

Let f : Z → Z be given by

f(z) = (

2z − 5, if z ≥ 0;

z + 5, if z < 0.

a) Is f an injective function?

b) Let u ∈ Z, u ≤ 5. show that u ∈ im(f).

c) Let v ∈ Z, v > 5, Show that v ∈ im(f) if and only if v + 5 is even .

[5,6,6]


1
Expert's answer
2021-10-18T16:38:10-0400

a) "2\cdot e^{2x-1}+5\cdot e^x-2=0"

"e^x=t" - new unknown

"\frac{2}{e}\cdot t^2+5\cdot t-2=0"

"D=25+\frac{16}{e}"

"t_1=\frac{-5-\sqrt{25+\frac{16}{e}}}{\frac{4}{e}}"

"t_2=\frac{-5+\sqrt{25+\frac{16}{e}}}{\frac{4}{e}}"

1)"e^x=t_1" - no solutions, because "t_1<0"

2) "e^x=t_2.\space x=ln\left( \frac{-5+\sqrt{25+\frac{16}{e}}}{\frac{4}{e}} \right)"

Answer: "x=ln\left( \frac{-5+\sqrt{25+\frac{16}{e}}}{\frac{4}{e}} \right)"

b)

"2^{4x}+2^{2x-1}>2\\
2^{2x}=t;\\
t^2+\frac{1}{2}t-2>0;"

"t_1=\frac{-1-\sqrt{33}}{4};"

"t_2=\frac{-1+\sqrt{33}}{4};"

"(t-t_1)(t-t_2)>0"

"(t<t_1)\vee(t>t_2)\\"

"t<t_1" -impossible, because "t_1<0"

"t>t_2;\\
2^{2x}>\frac{-1+\sqrt{33}}{4};\\
2x>log_2\left( \frac{-1+\sqrt{33}}{4} \right)\\
x>\frac{1}{2}\cdot log_2\left( \frac{-1+\sqrt{33}}{4} \right)\\
answer:\space x\in \left( \frac{1}{2}\cdot log_2\left( \frac{-1+\sqrt{33}}{4} \right).\infty \right)"

Question 2

"f(x)=\sqrt{2-x},\space dom(f)=(-\infty,2]"

f(x) is monotone increasing because "f_1(x)=2-x" is monotone decreasing and

"f_2(y)=\sqrt{y}" is monotone increasing , therefore "f(x)=f_2\left(f_1(x)\right)" is monotone decreasing as composition of monotone functions.

Monotine i funcrion f is bijection from dom(f) on im(f).

"f((-\infty,2])=[f(2), lim_{x\rarr -\infty}f(x)=[0.\infty)"

so im(f)="[0.\infty)"

Inverse function "g:R\rarr[-\infty,2)"

has dom(g)="[0,\infty)" and is determined by equation

f(y)=x, "or\space\sqrt{2-y}=x"

therefore

"y=2-x^2"

We have "g(x)=2-x^2"

Question 3

"f(z) = \begin{cases}
 2z-5 &{z\ge 0} \\
 z+5 &{z<0} 
\end{cases}"

a) Function f(z) is not injective because

f(0)=f(-10)=-5 and this is counterexample to injectivity

b) Let "u\in Z.u\le5"

For proving that "u\in im(f)" we must find solution for equation

f(z)=u

Let we search solution in region "(-\infty,0)"

In this case the last equation has form z+5=u;

From it we have z=u-5 and this solution correct if u<5

Now let us consider the case u=5, we have f(5)="2\cdot 5-5=5" and eqution

f(z)=u has solution anywhere if "u\le5" .

c) Let v ∈ Z, v > 5, Show that v ∈ im(f) if and only if v + 5 is even .

We consider two possibilities:

  1. z+5=v,z<0

But in this case z+5<5<v and this case impossible

2.

"2z-5=v,\\
z=\frac{v+5}{2}"

if v+5 is odd then this solution is not valid because "z\notin Z"

Contrary if v+5 is even, i.e. v+5=2k,"k\in Z,k>5"

we have z=k- is correct solution of equation f(z)=v, because z=k>5>0